OpenStudy (anonymous):

A chemist needs 90 milimeters of a 71% solution but has only 67% and 76% solutions avaliable. How many milimeters of each that should be mixed to get the desired solution?

OpenStudy (anonymous):

wow i did this exact same problem yesterday. where does it come from?

OpenStudy (anonymous):

a homework that I don't understand :(

OpenStudy (inkyvoyd):

Homework or test?

OpenStudy (anonymous):

Make up Test

OpenStudy (inkyvoyd):

A chemist needs 90 milimeters of a 71% solution but has only 67% and 76% 90 milimeters of a 71% -> 90*71% gives us the amount of the solute (the "salt" in salt water)

OpenStudy (inkyvoyd):

if x and y are the amount of 67% and 76% we add, then 90*71%=67%*x+76%*y

OpenStudy (inkyvoyd):

get rid of the percents (multiply by 100) 90*71=67x+76y

OpenStudy (inkyvoyd):

and, x+y=90, because the total amount of solution (the seawater) will be 90 ml

OpenStudy (anonymous):

no matter we can do it again. we need a variable so put \(x\)= number of ml of 76% solution, which will contribute \(.76x\) since the total is 90 ml the amount of 67% solution will be \(90-x\) and that will contribute \(.67(90-x)\) to the total. since you want 90 ml of 71% solution the total is \[.71\times 90=6.39\] and now we now have two expressions for the total so set them equal and solve \[.76x+.67(90-x)=6.39\] for \(x\)

OpenStudy (anonymous):

clear the decimals and get \[76x+67(90-x)=639\] and go from there

OpenStudy (anonymous):

oh, sorry what inkyvoid said